### 19 May 2022 v1

# Odd coloring of two subclasses of planar graphs thanks: Supported by the National Natural Science Foundation of China (No. 11871055). thanks: Mathematics Subject Classification (2020): 05C15, 05C10

Abstract
A proper coloring of a graph is odd if every non-isolated vertex has some color that appears an odd number of times on its neighborhood. Petruševski and Škrekovski conjectured in 2021 that every planar graph admits an odd $5$-coloring. We confirm this conjecture for outer-1-planar graphs and 2-boundary planar graphs, which are two subclasses of planar graphs.
Keywords: odd coloring; outer-1-planar graph; 2-boundary planar graph.
1 Introduction
All graphs in this paper are simple. A proper coloring of a graph is odd if every non-isolated vertex has some color that appears an odd number of times on its neighborhood. The odd chromatic number of $G$, denoted by $\chi_{o}(G)$, is the smallest number $k$ such that $G$ admits an odd $k$-coloring.
The notion of odd coloring was introduced by Petruševski and Škrekovski Petruvsevski2021ColoringsWN . They proved that every planar graph admits an odd $9$-coloring and conjectured the following.
Conjecture 1.
$\chi_{o}(G)\leq 5$ for every planar graph $G$.
It is easy to see that $\chi_{o}(C_{5})=5$, so the bound in this conjecture is sharp if it was correct. Recently, Petr and Portier Petr2022TheOC improved the bound for the odd chromatic number of planar graphs to 8, which is best known until now.
Towards Conjecture 1, it is natural to consider planar graphs with girth restriction. For each integer $t\leq 7$, let $g_{t}$ be the minimum integer such that every planar graph with girth at least $g_{t}$ has an odd $t$-coloring. Cranston Cranston2022OddCOO showed $g_{6}\leq 6$ and $g_{5}\leq 7$. Cho, Choi, Kwon, and Park Cho2022OddCO showed $g_{4}\leq 11$ and $g_{7}\leq 5$. Note that $g_{3}=+\infty$ as even cycles are not odd $3$-colorable.
The average degree $\mathrm{ad}(G)$ of a graph $G$ is $2|E(G)|/|V(G)|$. The maximum average degree $\mathrm{mad}(G)$ of a graph $G$ is $\max\{\mathrm{ad}(H)~|~H\subseteq G\}$. From the well-known Euler’s formula, one can conclude that $\mathrm{mad}(G)<2g/(g-2)$ for every planar graph $G$ with girth at least $g$. So a natural way to generate the above results on planar graphs with girth restriction is to investigate the same problem for graphs with bounded maximum average degree.
In view of this, Cranston Cranston2022OddCOO showed $\chi_{o}(G)\leq 5$ if $\mathrm{mad}(G)<20/7$, and $\chi_{o}(G)\leq 6$ if $\mathrm{mad}(G)<3$. Cho, Choi, Kwon, and Park Cho2022OddCO proved that if $t\geq 7$ and $\mathrm{mad}(G)<4t/(t-2)$ then $\chi_{o}(G)\leq t$, and if $G$ is a graph with $\mathrm{mad}(G)\leq 22/9$ containing no induced $5$-cycles then $\chi_{o}(G)\leq 4$. Note that the latter result implies $g_{4}\leq 11$ that was mentioned before.
An outerplanar graph is a graph that can be drawn in the plane without crossings so that all vertices belong to the unbounded face of the drawing. Caro, Petruševski and Škrekovski Caro2022RemarksOO proved that every outerplanar graph admits an odd 5-coloring, therefore, Conjecture 1 is valid when it is restricted to outerplanar graphs.
A graph is 1-planar if it can be drawn in the plane so that each edge is crossed at most once. Cranston, Lafferty, and Song Cranston2022ANO proved that $\chi_{o}(G)\leq 23$ for every 1-planar graph $G$.
The combination of 1-planarity and outerplanarity leads to outer-1-planar graphs, which are graphs admitting drawings in the plane in such a way that all vertices lie on the unbounded face and each edge crosses at most one another edge. For example, $K_{4}$ and $K_{2,3}$ are both outer-1-planar. Outer-1-planar graphs were first introduced by Eggleton MR846198 under the notion of outerplanar graphs with edge crossing number one MR846198 . They are also known as pseudo-outerplanar graphs MR2945171 ; MR3084275 ; MR3203677 in the literature. Note that every outer-1-planar graph is planar Hong2014ALA .
We introduce another subclass of planar graphs that generalizes outplanar graphs. A 2-boundary planar graph is a graph that has an embedding in the plane in such a way that all vertices lie on the boundary of at most two faces and no edge is crossed, and we call such an embedding 2-boundary plane graph. Figure 1 gives some examples of 2-boundary plane graphs. Note that they are not outerplanar, as we can find a $K_{4}$-minor in each graph.
Fig. 1: 2-boundary planar graphs
In this paper, we confirm Conjecture 1 for outer-1-planar graphs and 2-boundary planar graphs. This paper is organized as follows. In Section 2, we establish the structural properties of outer-1-planar graphs and 2-boundary planar graphs. We then introduce the notion of odd non-$k$-colorable minimal graphs in Section 3 and show that each configuration obtained in Section 2 is odd $k$-reducible for every integer $k\geq 5$. Combining the results of Sections 2 and 3, we prove in Section 4 that every outer-1-planar graph and every 2-boundary planar graph admits an odd 5-coloring. Note that this bound is sharp as $C_{5}$ is both outer-1-planar and 2-boundary planar and $\chi_{o}(C_{5})=5$.
2 Structural properties
The following structural lemma for outer-1-planar graphs is enough to be used later.
Lemma 2.
[dmtcs:8381, , Theorem 4, Li and Zhang] Every outer-$1$-planar graph $G$ contains
  1. (a2)
    a vertex $v$ of degree 11, or
  2. (b2)
    a vertex $v$ with $N_{G}(v)=\{u,w\}$ such that $uw\in E(G)$, or
  3. (c2)
    a vertex $v$ with $N_{G}(v)=\{u,w\}$ such that $uw\not\in E(G)$ and Gv+uwG-v+uw is outer-1-planar, or
  4. (d2)
    a cycle uxvyuuxvyu with $d_{G}(u)=d_{G}(v)=3$ and $uv\in E(G)$.
Hence in this section we mainly focus on the structure of 2-boundary planar graphs.
If all vertices of a 2-boundary plane graph $G$ lie on the boundaries of exactly two faces, then one of them is unbounded and the other is bounded. We call the unbounded face outside face and the bounded face inside face, denoted by $f_{out}$ and $f_{in}$ respectively. Other faces are called interior faces. By $V(f_{out})$ (resp. $V(f_{in})$) we denote the set of vertices on $f_{out}$ (resp. $f_{in}$). Note that it is possible that $V(f_{out})\cap V(f_{in})\neq\emptyset$ (see the last three pictures of Figure 1). If this occurs, we call vertices in $V(f_{out})\cap V(f_{in})$ shared vertices. Let
$\tilde{\Delta}(G)=\max\bigg\{d_{G}(u)~\big|~u\in V(G)\setminus\big(V(f_{out})\cap V(f_{in})\big)\bigg\}.$
For an edge $uv\in E(G)$ such that $\{u,v\}\subseteq V(f_{out})$ or $\{u,v\}\subseteq V(f_{in})$, it is a chordal edge if $u$ and $v$ are not consecutive on $f_{out}$ or $f_{in}$, respectively, and boundary edge otherwise. An edge $uv\in E(G)$ is an interconnected edge if $u\in V(f_{out})$, $v\in V(f_{in})$, and $\{u,v\}\cap V(f_{out})\cap V(f_{in})=\emptyset$.
A $k$-vertex is a vertex of degree $k$. A vertex with odd (resp. even) degree is an odd vertex (resp. even vertex). A neighbor of a vertex is odd neighbor if it is an odd vertex.
Lemma 3.
Every $2$-boundary planar graph $G$ contains one of the following configurations unless $G$ is isomorphic to $M$ and $P$ as described in Figure 2:
  1. (a3)
    a vertex $v$ of degree $1$;
  2. (b3)
    a triangle uvwuuvwu with $d_{G}(v)=2$;
  3. (c3)
    a path uvwuvw with $d_{G}(v)=2$ and $uw$ is a non-edge such that Gv+uwG-v+uw is $2$-boundary planar;
  4. (d3)
    a $3$-vertex $v$ with two odd neighbors;
  5. (e3)
    a triangle uvwuuvwu with $d_{G}(v)=d_{G}(w)=3$;
  6. (f3)
    two adjacent triangles xuvxxuvx and yuvyyuvy with $d_{G}(u)=d_{G}(v)=4,d_{G}(x)=3$ and $d_{G}(y)$ being odd;
  7. (g3)
    a triangle vxyvvxyv adjacent to two triangles uxvuuxvu and wvywwvyw such that $u\neq w$, $d_{G}(v)=d_{G}(x)=d_{G}(y)=4$, $d_{G}(u)=3$;
  8. (h3)
    a fan with center vertex $v$ and boundary path uxyzwuxyzw such that $d_{G}(v)=5$, $d_{G}(x)=d_{G}(z)=4$, and $d_{G}(y)=3$.
Fig. 2: 3-regular and 4-regular 2-boundary planar graphs
Suppose for a contradicting that $G$ is a $2$-boundary plane graph that contains none of the above configurations and has the minimum number of vertices.
If $G$ is an outerplane graph, then $G$ contains (a3), (b3), or (c3) (see zbMATH00790475 ; dmtcs:8381 ), a contradiciton. Hence $G$ has an outside face $f_{out}$ and an inside face $f_{in}$.
Clearly, $G$ is connected and $\delta(G)\geq 2$ by the absence of (a3).
Claim 1.
$\delta(G)\geq 3$.
If $v$ is a $2$-vertex in $G$, then $v$ is not a shared vertex as every shared vertex has degree at least 3. So we assume w.l.o.g. $v\in V(f_{out})\setminus V(f_{in})$. Let $N_{G}(v)=\{u,w\}$. Now $u,w\in V(f_{out})$. If $uw\in E(G)$, then (b3) occurs. If $uw\not\in E(G)$, then it is easy to see that Gv+uwG-v+uw is 2-boundary planar and thus (c3) occurs. Hence $\delta(G)\geq 3$. $\square$
Claim 2.
$G$ does not has any chordal edge.
If $G$ has a chordal edge, then choose one, say $uv$, such that there is no other chordal edge $xy$ such that $u,x,y,v$ lies in this ordering on $f_{out}$ (or $f_{in}$). This implies that there is a $2$-vertex $z$ such that zuvzzuvz forms a triangle, or there are two $2$-vertices $z_{1},z_{2}$ such that $uz_{1}z_{2}$ forms a path and $uz_{2}$ is a non-edge. The former case implies the appearance of (b3) and the later case implies the appearance of (c3). Hence $G$ does not has any chordal edge. $\square$
Claim 3.
If $v\in V(f_{out})\setminus V(f_{in})$ (resp. $v\in V(f_{in})\setminus V(f_{out})$) and $u_{1},u_{2},\ldots,u_{d-2}$ $(d\geq 4)$ are all neighbors of $v$ in $V(f_{in})\setminus V(f_{out})$ (resp. $V(f_{out})\setminus V(f_{in})$) lying in this ordering on $V(f_{in})$ (resp. $V(f_{out})$), then $u_{i}u_{i+1}$ is a boundary edge for each $1\leq i\leq d-3$, and moreover, $d_{G}(u_{i})=3$ for each $2\leq i\leq d-3$.
Suppose that there is some vertex $w$ such that $u_{i},w,u_{i+1}$ lie in this ordering on $f_{in}$ (resp. $f_{out}$) for some $1\leq i\leq d-3$. Since $G$ does not has any chordal edge by Claim 2, $w$ has at most two neighbors in $V(f_{in})$ (resp. $V(f_{out})$). Further, since $d_{G}(w)\geq 3$ by Claim 1, $w$ has at least one neighbor in $V(f_{out})$ (resp. $V(f_{in})$). Since no edge is crossed in $G$, the neighbor of $w$ in $V(f_{out})$ (resp. $V(f_{in})$) must be $v$. This is a contradiction as $w$ is not a neighbor of $v$.
So $u_{i}$ and $u_{i+1}$ are consecutive on $V(f_{in})$ (resp. $V(f_{out})$) for each $1\leq i\leq d-3$. If $u_{i}u_{i+1}\not\in E(G)$, then $u_{i}vu_{i+1}$ would be a path on $V(f_{in})$ (resp. $V(f_{out})$). This implies $v\in V(f_{in})$ (resp $v\in V(f_{out})$), a contradiction. Hence $u_{i}u_{i+1}\in E(G)$, i.e., $u_{i}u_{i+1}$ is a boundary edge. It gives $N_{G}(u_{i})=\{u_{i-1},u_{i+1},v\}$ for each $2\leq i\leq d-3$, implying $d_{G}(u_{i})=3$. $\square$
Now we divide the main proof into a seires of cases.
Case 1: $\tilde{\Delta}(G)\geq 6$.
Choose w.l.o.g. a vertex $v\in V(f_{out})\setminus V(f_{in})$ such that $d_{G}(v):=d\geq 6$. Since $G$ does not has any chordal edge, there are d2d-2 neighbors of $v$, say $u_{1},u_{2},\ldots,u_{d-2}$, in $V(f_{in})\setminus V(f_{out})$. We assume that $u_{1},u_{2},\ldots,u_{d-2}$ lie in this ordering on $f_{in}$. By Claim 3, $vu_{2}u_{3}v$ forms a triangle with $d_{G}(u_{2})=d_{G}(u_{3})=3$, and thus a configuration (e3).
Case 2: $\tilde{\Delta}(G)=5$.
Choose w.l.o.g. a vertex $v\in V(f_{out})\setminus V(f_{in})$ such that $d_{G}(v)=5$. Since $G$ does not has any chordal edge by Claim 2, there are three neighbors of $v$, say $u_{1},u_{2},u_{3}$, in $V(f_{in})\setminus V(f_{out})$. We assume that $u_{1},u_{2},u_{3}$ lie in this ordering on $V(f_{in})$. By Claim 3, $u_{1}u_{2}u_{3}$ is a path on $f_{in}$ and $d_{G}(u_{2})=3$. By the absence of the configuration (d3), $d_{G}(u_{1})$ and $d_{G}(u_{3})$ are even. Since $u_{1},u_{3}\not\in V(f_{out})$, they are not shared vertices and thus $d_{G}(u_{1})=d_{G}(u_{3})=4$ as $\tilde{\Delta}(G)=5$.
Since chordal edges are forbidden in $G$, for $i=1,3$, there is another vertex besides $v$, say $w_{i}$, acting as a neighbor of $u_{i}$ on $V(f_{out})\setminus V(f_{in})$. By Claim 3, $vw_{1}$ and $vw_{3}$ are both boundary edges on $f_{out}$. This implies the occurrence of the configuration (h3).
Case 3: $\tilde{\Delta}(G)=4$.
Choose w.l.o.g. a vertex $v\in V(f_{out})\setminus V(f_{in})$ such that $d_{G}(v)=4$. Since $G$ does not has any chordal edge by Claim 2, there are two neighbors of $v$, say $u_{1},u_{2}$, in $V(f_{in})\setminus V(f_{out})$. By Claim 3, $u_{1}u_{2}$ is a boundary edge on $f_{in}$. By the absence of the configuration (e3), $d_{G}(u_{1})\neq 3$ or $d_{G}(u_{2})\neq 3$. Since $u_{1},u_{2}\not\in V(f_{out})$, they are not shared vertices and thus $d_{G}(u_{1}),d_{G}(u_{2})\leq\tilde{\Delta}(G)$. It follows $d_{G}(u_{1})=4$ or $d_{G}(u_{2})=4$. We assume by symmetry that $d_{G}(u_{2})=4$. Since chordal edges are forbidden in $G$, there is another vertex besides $v$, say $w_{2}$, acting as a neighbor of $u_{2}$ on $V(f_{out})\setminus V(f_{in})$. By Claim 3, $vw_{2}$ is a boundary edge on $f_{out}$.
If $d_{G}(u_{1})=3$ and $d_{G}(w_{2})$ is odd, then the configuration (f3) appears.
If $d_{G}(u_{1})=3$ and $d_{G}(w_{2})$ is even, then $d_{G}(w_{2})=4$ as $w_{2}$ is not a shared vertex. So there is a vertex $u_{3}$ on $V(f_{in})\setminus V(f_{out})$ such that $w_{2}u_{3}\in E(G)$. By Claim 3, $u_{2}u_{3}$ is a boundary edge on $f_{in}$, and thus the configuration (g3) appears.
If $d_{G}(u_{1})=4$, then there is a vertex $w_{0}$ on $V(f_{out})\setminus V(f_{in})$ such that $w_{0}u_{1}\in E(G)$. By Claim 3, $w_{0}v$ is a boundary edge on $f_{out}$. If $d_{G}(w_{2})=3$, then the configuration (g3) appears. Hence $d_{G}(w_{2})=4$. So there is a vertex $u_{3}$ on $V(f_{in})\setminus V(f_{out})$ such that $w_{2}u_{3}\in E(G)$. By Claim 3, $u_{2}u_{3}$ is a boundary edge on $f_{in}$. If $d_{G}(u_{3})=3$, then the configuration (f3) appears. Hence $d_{G}(u_{3})=4$.
Let $w_{1}:=v$. For each $i\geq 1$, we repeat the above arguments via letting $w_{i+1},u_{i+1},u_{i+2}$ play the roles of $w_{i},u_{i},u_{i+1}$, respectively, and would find two non-shared $4$-vertices $w_{i+2}$ and $u_{i+3}$ such that $w_{i+2}u_{i+2},w_{i+2}u_{i+3}\in E(G)$ and $w_{i+1}w_{i+2},u_{i+2}u_{i+3}$ are boundary edges unless (f3) or (g3) occurs. Since the graph $G$ is finite, if we forbid the configurations (f3) and (g3) then this iterative process would finally end without finding any new vertex only if $G\cong P$.
Case 4: $\tilde{\Delta}(G)=3$.
Suppose that $u$ is a shared vertex. Let $v\in V(f_{out})$ and $w\in V(f_{in})$ such that $uv,uw$ are boundary edges and $v,w$ lie on the boundary of a same interior face. If $v$ is a shared vertex, then $uv$ is a chordal edge as otherwise $uv$ is a multi-edge. Hence $v$ is not a shared vertex and similarly $w$ is not either. This gives $d_{G}(v)=d_{G}(w)=\tilde{\Delta}(G)=3$ as $\delta(G)\geq 3$. Moreover, it is easy to see that the unique neighbor of $v$ on $f_{in}$ is exactly $w$ as otherwise $w$ can only has degree two. However, the triangle uvwuuvwu forms the configuration (e3), which is forbidden.
Therefore, $G$ has no shared vertex and thus $\Delta(G)=\tilde{\Delta}(G)=3$. Since $\delta(G)\geq 3$ by Claim 1, $G$ is 3-regular. Since $G$ does not has any chordal edge by Claim 2, $G\cong M$. ∎
3 Reducibilities
Let $\mathcal{G}$ be a class of graphs $G$ such that $H\in\mathcal{G}$ for every subgraph $H$ of $G$. An odd non-$k$-colorable minimal graph (w.r.t. $\mathcal{G}$) is a graph $G\in\mathcal{G}$ such that
  • $\chi_{o}(G)>k$ and
  • $\chi_{o}(H)\leq k$ for every graph $H\in\mathcal{G}$ with $|V(H)|+|E(H)|<|V(G)|+|E(G)|$.
A configuration is odd $k$-reducible in $\mathcal{G}$ if it cannot occur in an odd non-$k$-colorable minimal graph (w.r.t. $\mathcal{G}$).
For an odd $k$-coloring $\tau$ of some subgraph $H$ of $G$ and for a vertex $v\in V(H)$, let $\tau_{o}(v)$ denote the unique color that appears an odd number of times on $N_{H}(v)$ if such a color exists; otherwise, we set $\tau_{o}(v)=0$. Note that $\tau_{o}(v)\not=0$ implies $d_{G}(v)$ is odd.
Define $\tau(N_{H}(v))=\{\tau(u)~|~u\in N_{H}(v)\}$ and $\tau_{o}(N_{H}(v))=\{\tau_{o}(u)~|~u\in N_{H}(v)\}$. While extending $\tau$ to an odd $k$-coloring of $G$, the set of colors that cannot be used by a being colored vertex $v$ is therefore
$F^{\tau}(v)=[k]\cap\bigg(\tau(N_{H}(v))\cup\tau_{o}(N_{H}(v))\bigg).$
Formally, the set $F^{\tau}(v)$ is called the forbidden set for $v$, and the set
$A^{\tau}(v)=[k]\setminus F^{\tau}(v)$
is called the available set for $v$, where $[k]$ denotes the set of colors $\{1,2,\ldots,k\}$.
Note that we shall always check the oddness of $v$ (i.e. whether there is some color that appears an odd number of times on the neighborhood of $v$) after the extension.
Lemma 4.
A vertex $v$ of degree 1 is odd $k$-reducible in $\mathcal{G}$ for every integer $k\geq 3$.
If $G$ is an odd non-$k$-colorable minimal graph containing such a configuration, then $G^{\prime}=G-v$ is odd $k$-colorable. Let $\tau$ be an odd $k$-coloring of $G^{\prime}$ and let $u$ be the neighbor of $v$. We extend $\tau$ to an odd $k$-coloring of $G$ by coloring $v$ with a color in $k]\setminus F^{\tau}(v)$. This is possible since $F^{\tau}(v)=[k]\cap\{\tau(u),\tau_{o}(u)\}$. ∎
Lemma 5.
A triangle uvwuuvwu with $d_{G}(v)=2$ is odd $k$-reducible in $\mathcal{G}$ for every integer $k\geq 5$.
If $G$ is an odd non-$k$-colorable minimal graph containing such a configuration, then $G^{\prime}=G-v$ is odd $k$-colorable. Let $\tau$ be an odd $k$-coloring of $G^{\prime}$. We extend $\tau$ to an odd $k$-coloring of $G$ by coloring $v$ with a color in $k]\setminus F^{\tau}(v)$. This is possible since $F^{\tau}(v)=[k]\cap\{\tau(u),\tau(w),\tau_{o}(u),\tau_{o}(w)\}$. ∎
Lemma 6.
A path uvwuvw with $d_{G}(v)=2$ and $uw$ being a non-edge is odd $k$-reducible in $\mathcal{G}$ for every integer $k\geq 5$ if $G-v+uw\in\mathcal{G}$.
If $G$ is an odd non-$k$-colorable minimal graph containing such a configuration, then $G^{\prime}=G-v+uw$ is odd $k$-colorable as $G^{\prime}\in\mathcal{G}$. Let $\tau$ be an odd $k$-coloring of $G^{\prime}$. Clearly, $\tau(u)\neq\tau(w)$. Assume w.l.o.g. that $\tau(u)=1$ and $\tau(w)=2$. Note that $N_{G^{\prime}}(u)\setminus\{w\}=N_{G}(w)\setminus\{v\}$ and $N_{G^{\prime}}(w)\setminus\{u\}=N_{G}(w)\setminus\{v\}$. We divide the proofs into three cases by symmetry.
Case 1. $\tau_{o}(u)=2$ and $\tau_{o}(w)=1$
Now every color in $\{3,4,\ldots,k\}$ appears an even number of times on both $N_{G}(u)\setminus\{v\}$ and $N_{G}(w)\setminus\{v\}$. Hence we extend $\tau$ to an odd $k$-coloring of $G$ by coloring $v$ with $3$.
Case 2. $\tau_{o}(u)=2$ and $\tau_{o}(w)\not=1$
Similarly, every color in $\{3,4,\ldots,k\}$ appears an even number of times on $N_{G}(u)\setminus\{v\}$. If $\tau_{o}(w)=0$, then there is color different from 1, say 3, appearing an odd number of times on $N_{G}(w)\setminus\{v\}$. If $\tau_{o}(w)\neq 0$, then assume w.l.o.g. that $\tau_{o}(w)=3$. In each case we are possible to extend $\tau$ to an odd $k$-coloring of $G$ by coloring $v$ with $4$.
Case 3. $\tau_{o}(u)\not=2$ and $\tau_{o}(w)\not=1$
We restrict $\tau$ to a $k$-coloring $\phi$ of GvG-v. If $\tau_{o}(u)\not=0$ (assume w.l.o.g. $\tau_{o}(u)=3$), then the colors $2$ and $3$ appear an odd number of times on $N_{G}(u)\setminus\{v\}$. If $\tau_{o}(u)=0$, then there is a color different from 1 and 2 that appears an odd number of times on $N_{G}(u)\setminus\{v\}$. Hence in each case we find a color $a\neq 1,2$ that appears an odd number of times on $N_{G}(u)\setminus\{v\}$. Similarly, we can find a color $b\neq 1,2$ that appears an odd number of times on $N_{G}(w)\setminus\{v\}$. So we extend $\phi$ to an odd $k$-coloring of $G$ by coloring $v$ with a color in $\{3,4,\ldots,k\}\setminus\{a,b\}$. ∎
Lemma 7.
A $3$-vertex $v$ with two odd neighbors is odd $k$-reducible in $\mathcal{G}$ for every integer $k\geq 5$.
If $G$ is an odd non-$k$-colorable minimal graph containing a $3$-vertex $v$ with $N_{G}(v)=\{x,y,z\}$ such that $x$ and $y$ are odd vertices of $G$, then $G^{\prime}=G-v$ has an odd $k$-coloring $\tau$. Now $F^{\tau}(v)=[k]\cap\{\tau(x),\tau(y),\tau(z),\tau_{o}(x),\tau_{o}(y),\tau_{o}(z)\}$. Since $x$ and $y$ are even vertices of $G^{\prime}$, $\tau_{o}(x)=\tau_{o}(y)=0$. It follows $|A^{\tau}(v)|\geq k-4\geq 1$. Hence we can extend $\tau$ to an odd $k$-coloring of $G$ by coloring $v$ with a color in $A^{\tau}(v)$. Note that the oddness of $v$ can be always guaranteed as $v$ is an odd vertex of $G$. ∎
Lemma 8.
A triangle uvwuuvwu with $d_{G}(v)=d_{G}(w)=3$ is odd $k$-reducible in $\mathcal{G}$ for every integer $k\geq 5$.
If $G$ is an odd non-$k$-colorable minimal graph containing such a configuration, then $G^{\prime}=G-\{v,w\}$ has an odd $k$-coloring $\tau$. Let $x\in N_{G}(v)\setminus\{u,w\}$ and $y\in N_{G}(w)\setminus\{u,v\}$. Note that it may be possible that $x=y$.
If $d_{G}(u)$ is odd, then such a configuration is odd $k$-reducible in $\mathcal{G}$ for every integer $k\geq 5$ by Lemma 7. Hence we assume that $d_{G}(u)$ is even. It follows that $d_{G^{\prime}}(u)$ is even and thus $\tau_{o}(u)=0$. So $A^{\tau}(v)=[k]\setminus\{\tau(x),\tau_{o}(x),\tau(u)\}$ and $|A^{\tau}(v)|\geq k-3\geq 2$. We assume w.l.o.g. $A^{\tau}(v)=\{1,2\}$.
We extend $\tau$ to an odd $k$-coloring $\phi$ of $G^{\prime\prime}=G-w$ by coloring $v$ with 1. Since $d_{G^{\prime\prime}}(v)=2$, $\phi_{o}(v)=0$. It follows that $A^{\phi}(w)=[k]\setminus\{\phi(y),\phi_{o}(y),\phi(u),\phi_{o}(u),\phi(v)\}$. If $|A^{\phi}(w)|\geq 1$, then we extend $\phi$ to an odd $k$-coloring of $G$ by coloring $w$ with a color in $A^{\phi}(w)$. Hence we assume that $k=5$ and $\{\phi(y),\phi_{o}(y),\phi(u),\phi_{o}(u)\}=\{2,3,4,5\}$.
In this situation we adjust the coloring $\phi$ to $\phi^{\prime}$ by recoloring $v$ with 2. This is realizable as $A^{\tau}(v)=\{1,2\}$. If $x\neq y$, then $\{\phi^{\prime}(y),\phi^{\prime}_{o}(y),\phi^{\prime}(u),\phi^{\prime}_{o}(u),\phi^{\prime}(v)\}=\{\phi(y),\phi_{o}(y),\phi(u),\phi_{o}(u),\phi(v)\}=\{2,3,4,5\}$ and $1\in A^{\phi}(w)$. If $x=y$, then $2\notin\{\tau(x),\tau_{o}(x)\}=\{\phi(y),\phi_{o}(y)\}$. It follows $\phi^{\prime}_{o}(y)=\phi_{o}(y)$ and thus $\{\phi^{\prime}(y),\phi^{\prime}_{o}(y),\phi^{\prime}(u),\phi^{\prime}_{o}(u),\phi^{\prime}(v)\}=\{\phi(y),\phi_{o}(y),\phi(u),\phi_{o}(u),2\}=\{2,3,4,5\}$ and $1\in A^{\phi}(w)$. Hence in each case we complete an odd $5$-coloring of $G$ by coloring $w$ with 1. ∎
Lemma 9.
Two adjacent triangles xuvxxuvx and yuvyyuvy with $d_{G}(u)=d_{G}(v)=4,d_{G}(x)=3$ and $d_{G}(y)$ being odd is odd $k$-reducible in $\mathcal{G}$ for every integer $k\geq 5$.
If $G$ is an odd non-$k$-colorable minimal graph containing such a configuration, then $G^{\prime}=G-x$ is odd $k$-colorable. Let $\tau$ be an odd $k$-coloring of $G^{\prime}$. Let $x^{\prime}\in N_{G}(x)\setminus\{u,v\}$, $u^{\prime}\in N_{G}(u)\setminus\{x,y,v\}$, and $v^{\prime}\in N_{G}(v)\setminus\{x,y,u\}$. Note that $\tau_{o}(u)=\tau_{o}(v)=0$.
If we are able to color $x$ with a color $a\in[k]\setminus\{\tau(u),\tau(v),\tau(y),\tau(x^{\prime}),\tau_{o}(x^{\prime})\}$, then we obtain an odd $k$-coloring of $G$. Note that the oddnesses on $u$ and $v$ can be guaranteed as $|\{a,\tau(v),\tau(y)\}|=3$ and $|\{a,\tau(u),\tau(y)\}|=3$.
Hence we assume below that $k=5$ and $\{\tau(u),\tau(v),\tau(y),\tau(x^{\prime}),\tau_{o}(x^{\prime})\}=[5]$. Assume w.l.o.g. $\tau(u)=1,\tau(v)=2,\tau(y)=3,\tau(x^{\prime})=4,\tau_{o}(x^{\prime})=5$. If $\tau(u^{\prime})\neq 2$ and $\tau(v^{\prime})\neq 1$, then we complete an odd $5$-coloring of $G$ by coloring $x$ with 3. Otherwise, we assume w.l.o.g. $\tau(u^{\prime})=2$.
Now erase the color on $u$ and denote the resulting odd $k$-coloring of $G^{\prime\prime}=G^{\prime}-u$ by $\phi$. Since $d_{G^{\prime\prime}}(y)$ is even, $\phi_{o}(y)=0$. Since $d_{G^{\prime\prime}}(v)=2$, $\phi_{o}(v)=0$. Hence $F^{\phi}(u)=[5]\cap\{\phi(y),\phi(v),\phi(u^{\prime}),\phi_{o}(u^{\prime})\}=[5]\cap\{2,3,\phi_{o}(u^{\prime})\}$ and $A^{\phi}(u)=\{1,4,5\}\setminus\{\phi_{o}(u^{\prime})\}$. We color $u$ with a color $a\in A^{\phi}(u)\setminus\{1\}$ and then complete an odd 5-coloring of $G$ by coloring $x$ with 1. ∎
Lemma 10.
A triangle vxyvvxyv adjacent to two triangles uxvuuxvu and wvywwvyw such that $u\neq w$, $d_{G}(v)=d_{G}(x)=d_{G}(y)=4$, and $d_{G}(u)=3$ is odd $k$-reducible in $\mathcal{G}$ for every integer $k\geq 5$.
If $G$ is an odd non-$k$-colorable minimal graph containing such a configuration, then $G^{\prime}=G-u$ has an odd $k$-coloring $\tau$. Let $x^{\prime}\in N_{G}(x)\setminus\{u,v,y\}$, $y^{\prime}\in N_{G}(y)\setminus\{x,v,w\}$, and $u^{\prime}\in N_{G}(u)\setminus\{x,v\}$.
Now $F^{\tau}(u)=[k]\cap\{\tau(x),\tau(v),\tau(u^{\prime}),\tau_{o}(x),\tau_{o}(v),\tau_{o}(u^{\prime})\}$. If $k\geq 7$, or $k=6$ and $|F^{\tau}(u)|\leq 5$, or $k=5$ and $|F^{\tau}(u)|\leq 4$, then we extend $\tau$ to an odd $k$-coloring of $G$ by coloring $u$ with a color in $k]\setminus F^{\tau}(u)$. If $k=6$ and $|F^{\tau}(u)|=6$, then assume w.l.o.g. $\tau_{o}(x)=1$ and $\tau_{o}(v)=2$. However $\tau_{o}(x)=1$ implies $\tau(y)=1$ but $\tau_{o}(v)=2$ implies $\tau(y)=2$, a contradiction. Hence we assume $k=5$ and $5]\cap\{\tau(x),\tau(v),\tau(u^{\prime}),\tau_{o}(u^{\prime}),\tau_{o}(x),\tau_{o}(v)\}=[5]$ below. It follows that $\tau_{o}(x)\neq 0$ or $\tau_{o}(v)\neq 0$.
Case 1. $\tau_{o}(x)\neq 0$.
Assume w.l.o.g. $\tau(x)=1$ and $\tau_{o}(x)=2$. It follows $\tau(y)=2$ and $\tau(v)=\tau(x^{\prime})$. Assume $\tau(v)=3$.
Subcase 1.1 $\tau_{o}(v)\neq 0$.
This situation implies $\tau_{o}(v)=2$ and $\tau(w)=1$. Since $5]=\{\tau(x),\tau(v),\tau(u^{\prime}),\tau_{o}(x),\tau_{o}(v),\tau_{o}(u^{\prime})\}=\{1,2,3,\tau(u^{\prime}),\tau_{o}(u^{\prime})\}$, $\{\tau(u^{\prime}),\tau_{o}(u^{\prime})\}=\{4,5\}$. Recolor $x$ with a color in $\{4,5\}\setminus\{\tau_{o}(x^{\prime})\}$ and then color $u$ with 1. One can check that the resulting coloring is an odd $5$-coloring of $G$.
Subcase 1.2 $\tau_{o}(v)=0$.
This situation implies $|\{\tau(w),\tau(x),\tau(y)\}|=3$. So we assume w.l.o.g. $\tau(w)=4$. Since $5]=[5]\cap\{\tau(x),\tau(v),\tau(u^{\prime}),\tau_{o}(x),\tau_{o}(v),\tau_{o}(u^{\prime})\}=\{1,2,3,\tau(u^{\prime}),\tau_{o}(u^{\prime})\}$, $\{\tau(u^{\prime}),\tau_{o}(u^{\prime})\}=\{4,5\}$.
If $\tau_{o}(x^{\prime})=4$, then recolor $x$ with 5 and color $u$ with 1. If $\tau_{o}(x^{\prime})\not=4$ and $\tau(y^{\prime})\not=3$, then recolor $x$ with 4 and color $u$ with 1. If $\tau_{o}(x^{\prime})\not=4$ and $\tau(y^{\prime})=3$, then recolor $x$ with 4 and $v$ with a color in $\{1,5\}\setminus\{\tau_{o}(w)\}$, and color $u$ with 3. In each case we obtain an odd $5$-coloring of $G$.
Case 2. $\tau_{o}(v)\neq 0$.
Assume w.l.o.g. $\tau(v)=1$ and $\tau_{o}(v)=2$. It follows $\tau(y)=2$ and $\tau(w)=\tau(x)$. Assume $\tau(x)=3$.
Subcase 2.1 $\tau_{o}(x)\neq 0$.
This situation implies $\{\tau_{o}(x),\tau(x^{\prime})\}=\{1,2\}$. Since $5]=\{\tau(x),\tau(v),\tau(u^{\prime}),\tau_{o}(x),\tau_{o}(v),\tau_{o}(u^{\prime})\}=\{1,2,3,\tau(u^{\prime}),\tau_{o}(u^{\prime})\}$, $\{\tau(u^{\prime}),\tau_{o}(u^{\prime})\}=\{4,5\}$. Recolor $x$ with a color in $\{4,5\}\setminus\{\tau_{o}(x^{\prime})\}$ and then color $u$ with 3. One can check that the resulting coloring is an odd $5$-coloring of $G$.
Subcase 2.2 $\tau_{o}(x)=0$.
This situation implies $|\{\tau(x^{\prime}),\tau(y),\tau(v)\}|=3$. So we assume w.l.o.g. $\tau(x^{\prime})=4$. Since $5]=[5]\cap\{\tau(x),\tau(v),\tau(u^{\prime}),\tau_{o}(x),\tau_{o}(v),\tau_{o}(u^{\prime})\}=\{1,2,3,\tau(u^{\prime}),\tau_{o}(u^{\prime})\}$, $\{\tau(u^{\prime}),\tau_{o}(u^{\prime})\}=\{4,5\}$. If $\tau_{o}(x^{\prime})\neq 5$, then recolor $x$ with 5 and color $u$ with 3. If $\tau_{o}(x^{\prime})=5$, then recolor $x$ with 1 and color $v$ with a color in $\{4,5\}\setminus\{\tau_{o}(w)\}$, and color $u$ with 3. In each case we obtain an odd $5$-coloring of $G$. ∎
Lemma 11.
A fan with center vertex $v$ and boundary path uxyzwuxyzw such that $d_{G}(v)=5$, $d_{G}(x)=d_{G}(z)=4$, and $d_{G}(y)=3$ is odd $k$-reducible in $\mathcal{G}$ for every integer $k\geq 5$.
If $G$ is an odd non-$k$-colorable minimal graph containing such a configuration, then $G^{\prime}=G-y$ has an odd $k$-coloring $\tau$. Let $x^{\prime}\in N_{G}(x)\setminus\{u,v,y\}$ and $z^{\prime}\in N_{G}(z)\setminus\{y,v,w\}$. Note that $\tau_{o}(v)=0$ as $d_{G^{\prime}}(v)=4$.
Now $F^{\tau}(y)=[k]\cap\{\tau(x),\tau(v),\tau(z),\tau_{o}(x),\tau_{o}(z)\}$. If $k\geq 6$, then we color $y$ with a color in $k]\setminus F^{\tau}(y)$ to complete an odd $k$-coloring of $G$. So we assume $k=5$ below. Assume w.l.o.g. $\tau(x)=1,\tau(v)=2,\tau(z)=3,\tau_{o}(x)=4,\tau_{o}(z)=5$. It follows $\tau(x^{\prime})=\tau(z^{\prime})=2,\tau(u)=4,$ and $\tau(w)=5$.
If $\{\tau_{o}(x^{\prime}),\tau_{o}(u)\}\neq\{3,5\}$, then recolor $x$ with a color in $\{3,5\}\setminus\{\tau_{o}(x^{\prime}),\tau_{o}(u)\}$ and color $y$ with 1. If $\{\tau_{o}(z^{\prime}),\tau_{o}(w)\}\neq\{1,4\}$, then recolor $z$ with a color in $\{1,4\}\setminus\{\tau_{o}(z^{\prime}),\tau_{o}(w)\}$ and color $y$ with 3. In each case one can easily check that the resulting coloring is an odd $5$-coloring of $G$.
If $\{\tau_{o}(x^{\prime}),\tau_{o}(u)\}=\{3,5\}$ and $\{\tau_{o}(z^{\prime}),\tau_{o}(w)\}=\{1,4\}$, then the color 2 appears an even number of times (at least twice) on $N_{G^{\prime}}(u)$ and $N_{G^{\prime}}(w)$. We recolor $x$ with $\tau_{o}(u)$ and change the color of $v$ from $2$ to $1$. Note that $\tau_{o}(u)\neq\tau_{o}(x^{\prime})$ and the color 2 now appears an odd number of times on $N_{G}(u)$ and $N_{G}(w)$ no matter which color is assigned to $y$. Hence we can finally complete an odd $5$-coloring of $G$ by coloring $y$ with 2. ∎
4 Odd 5-colorabilities
Theorem 12.
Every outer-1-planar graph admits an odd 5-coloring.
Suppose for a contradiction that there is an odd non-5-colorable minimal outer-1-planar graph $G$. By Lemma 2, $G$ contains at least one configuration among (a2), (b2), (c2), and (d2). However, (a2), (b2), (c2), and (d2) are respectively odd 5-reducible by Lemma 4, 5, 6, and 8, a contradiction. ∎
Theorem 13.
Every 2-boundary planar graph admits an odd 5-coloring.
Suppose for a contradiction that there is an odd non-$5$-colorable minimal 2-boundary planar graph $G$.
If $G\cong M$, then $\chi(G)\leq 3$ by the classic Brooks’ theorem on vertex coloring. Since every vertex of $G$ has odd degree, every proper coloring of $G$ is an odd coloring. Hence $\chi_{o}(G)\leq\chi(G)\leq 3$, a contradiction.
If $G\cong P$, then assume $V(f_{out})=\{u_{1},u_{2},\dots,u_{n}\}$, $V(f_{in})=\{v_{1},v_{2},\dots,v_{n}\}$, $N_{G}(v_{1})=\{u_{n},u_{1},v_{n},v_{2}\}$, $N_{G}(v_{n})=\{u_{n-1},u_{n},v_{n-1},v_{1}\}$, and $N_{G}(v_{i})=\{u_{i-1},u_{i},v_{i-1},v_{i+1}\}$ for each $i=2,3,\dots,n-1$. If $n$ is even, we construct an odd $4$-coloring $\tau$ of $G$ by letting $\tau(u_{i})=1,\tau(u_{i+1})=2,\tau(v_{i})=3$, and $\tau(v_{i+1})=4$ for $i=1,3,\dots,n-1$. If $n$ is odd, we construct an odd $5$-coloring $\tau$ of $G$ by letting $\tau(u_{i})=4,\tau(u_{i+1})=5$ for $i=1,3,\ldots,n-4,n-1$, $\tau(u_{n-2})=3$, $\tau(v_{i})=1$, $\tau(v_{i+1})=2$ for $i=1,3,\dots,n-2$, and $\tau(v_{n})=3$. Hence $\chi_{o}(G)\leq 5$, a contradiction.
It follows $G\not\cong M$ and $G\not\cong P$ and thus $G$ contains at least one configuration among (a3), (b3), (c3), (d3), (e3), (f3), (g3), and (h3) by Lemma 3. However, (a3), (b3), (c3), (d3), (e3), (f3), (g3), and (h3) are respectively odd 5-reducible by Lemma 4, 5, 6, 7, 8, 9, 10, and $1$, a contradiction.∎